OpenStudy (blackops2_143):

A card was selected at random from a standard deck of cards. The suit of the card was recorded, and then the card was put back in the deck. The table shows the results after 40 trials. What is the relative frequency of selecting a heart? Outcome: Club Diamond Heart Spade number of trials: 8 12 14 6 A. 15% B. 25% C. 27% D. 35%

OpenStudy (ineedhelplz):

Relative Frequency: Total number of Diamonds / Total number of Cards Drawn

OpenStudy (ineedhelplz):

So, 12/40 = ?

OpenStudy (blackops2_143):

0.3

OpenStudy (ineedhelplz):

Right, now last step: 0.3 * 100 = ?

OpenStudy (blackops2_143):

30

OpenStudy (ineedhelplz):

Oh wait this is the relative frequency for diamond so sorry. Try again with heart.

OpenStudy (ineedhelplz):

So, 14/40 = ?

OpenStudy (ineedhelplz):

Sorry :(

OpenStudy (blackops2_143):

0.35

OpenStudy (ineedhelplz):

Then multiply that by 100

OpenStudy (blackops2_143):

35

OpenStudy (ineedhelplz):

Yes 35% So, the answer is D!

OpenStudy (ineedhelplz):

Sorry for messing up :/

OpenStudy (blackops2_143):

it's ok thanks so much for the help!